OpenStudy (anonymous):

the sum of the digits of a two digit number is 7. When the digits are reversed , the number is increased by27. Find the original number..

OpenStudy (anonymous):

solve: a+b=7 10a+b=27+10b+a

OpenStudy (anonymous):

10a+b=27+10b+a 9a=27+9b a=3+b a+b=7 a=7-b a=3+b 7-b=3+b 2b=10 b=5 a=7-b a=7-5 a=2 So the numbers are: 10a+b=20+5=25 a+10b=2+50=52 The original number is 25

OpenStudy (anonymous):

I agree with fiddlearound

OpenStudy (anonymous):

Do you understand what he did??

OpenStudy (anonymous):

i'll try